OpenStudy (anonymous):

How to do a modular arithmetic on a calculator? is there a way? I'm using casio fx-95MS

jimthompson5910 (jim_thompson5910):

What example problem do you want to work with?

jimthompson5910 (jim_thompson5910):

or I could make one up if you want?

OpenStudy (anonymous):

for example; 15^5 mod 128

jimthompson5910 (jim_thompson5910):

well 15^5 = 759,375 divide this by 128 (759,375)/128 = 5,932.6171875 We don't get a whole number as a result, but we can pull out the whole number part which is 5,932 The fractional or decimal part is 0.6171875 but we'll ignore it. ------------------------------------------------------- Now multiply the whole part 5,932 by the modulus 128 to get: 5,932*128 = 759,296 Then subtract the value of 15^5 which was 759,375 and 759,296 so, 759,375 - 759,296 = 79 So in the end, the remainder is 79 which means 759,375 mod 128 = 79 and 15^5 mod 128 = 79

jimthompson5910 (jim_thompson5910):

that example is a bit big, here's a simpler example 60 mod 7 = 4 but how did I find this? ------------------------------------------------------ Step 1) Divide 60 by the modulus 7: 60/7 = 8.57142857142858 ------------------ Step 2) Pull out the whole part whole part is 8 ignore the decimal part ------------------ step 3) Multiply the whole part (8) by the modulus (7) 8*7 = 56 ------------------ Step 4) Subtract the previous result from the original number 60 - 56 = 4 The remainder is 4. Therefore, 60 mod 7 = 4

OpenStudy (anonymous):

ohh my thank you so mucchh!! all clear to me now! :D

jimthompson5910 (jim_thompson5910):

I'm glad it is